A car accelerates with the acceleration a(s) = k · s^n m/s^2, where s is distance in meters, k = 1.8 and n = 0.725. (Unless otherwise stated, assume that the car has a  negligible initial speed, which only means that the answer does not become trivial  and that the car starts at s = 0.)
How far has the car driven when it has reached a speed of 1.4 m/s?
1
Expert's answer
2021-05-30T15:49:19-0400

"v^2=v_o^2+2as"

"v^2=2as"

"1.4^2=2*k*s^n*s"

"1.4^2=2\cdot \:1.8ss^{0.725}"

"1.4^2=3.6s^{0.725}s"

"\mathrm{Use\:the\:following\:exponent\:property}:\quad a^n=\left(a^{\frac{1}{m}}\right)^{\left(n\cdot \:m\right)}"

"1.4^2=3.6u^{29}u^{40}"

"s=\left(\frac{1.96}{3.6}\right)^{\frac{40}{69}}"

"1.4^2=3.6\left(s^{\frac{1}{40}}\right)^{29}\left(s^{\frac{1}{40}}\right)^{40}"

"s=\left(\frac{1.96}{3.6}\right)^{\frac{40}{69}}=s=0.70295 m"